Circular No. 7445

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of nine additional Kreutz sungrazing comets detected by SOHO with the C2 coronagraph, with measurements by D. Hammer (the reductions and orbital elements by B. G. Marsden appearing on the MPECs specified below). Comets C/1998 J3 and C/1998 J4 were discovered by Hammer in a search of archival data. The remaining comets were found via the SOHO website by M. Oates (C/1998 K12, C/1998 K13, C/1998 K14, C/1998 K15); M. Boschat (C/2000 L4, C/2000 L5, C/2000 M1); M. Meyer and X. Leprette (C/2000 L4); J. Shanklin (C/2000 L5); and P. Shrekby (C/2000 M1).

1998 UT  α_{2000}  δ_{2000}  MPEC  Comet
May 4.666 2^{h}50.^{m}3 +14^{o}43′ 2000-M27  C/1998 J3
6.227 2 56.3 +15 09  2000-M28  C/1998 J4
20.967 3 51.6 +18 34  2000-M29  C/1998 K12
21.769 3 51.9 +18 30  2000-M30  C/1998 K13
26.424 4 12.7 +19 31  2000-M31  C/1998 K14
28.584 4 19.7 +19 46  2000-M32  C/1998 K15
2000 UT  α_{2000}  δ_{2000}  MPEC  Comet
June 14.576 5^{h}29.^{m}0 +21^{o}32′ 2000-M33  C/2000 L4
15.576 5 32.0 +21 41  2000-M34  C/2000 L5
17.757 5 42.1 +21 34  2000-M35  C/2000 M1

THE EDGAR WILSON AWARD With reference to IAUC 6936 (and also to IAUC 7223), the Smithso-

nian Astrophysical Observatory wishes to announce that the 2000 Edgar Wilson Award for the discovery of comets by amateur astronomers is being divided among the following three individuals or groups: Daniel W. Lynn, Kinglake West, Victoria, Australia, for C/1999 N2; Korado Korlević, Višnjan, Croatia, for P/1999 WJ_{7}; and Gary Hug and Graham E. Bell, Eskridge, KS, U.S.A., for P/1999 X1.

SUPERNOVA 2000bs IN UGC 10710 CCD R magnitude estimates by K. Hornoch, Lelekovice, Czech Re-

public: May 5.954 UT, 17.0; 14.900, 17.2; June 1.957, 17.3; 17.973, 17.2.

2000 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green